graph the expression 1-tan^2x/1+tan^2x, and make a conjecture about another expression that is equivalent to another.


1
Expert's answer
2021-10-21T10:19:07-0400

Let us graph the function "y=\frac{1-\tan^2x}{1+\tan^2x}:"



It follows from the graph that "y=\cos(2x)."

Indeed, "\frac{1-\tan^2x}{1+\tan^2x}
=\frac{1-\tan^2x}{\frac{1}{\cos^2 x}}
=\cos^2x(1-\frac{\sin^2x}{\cos^2x})
=\cos^2x-\sin^2 x=\cos(2x)."